# Qudit Clauser-Horne-Shimony-Holt Inequality and Nonlocality from Wigner Negativity

### v4

@@ Abstract
Nonlocality is an essential concept that distinguishes quantum from classical models and has been extensively studied in systems of qubits. For higher-dimensional systems, certain results for their two-level counterpart, like Bell violations with stabilizer states and Clifford operators, do not generalize. On the other hand, similar to continuous variable systems, Wigner negativity is necessary for nonlocality in qudit systems. We propose a new generalization of the CHSH inequality for qudits by inquiring correlations related to the Wigner negativity of stabilizer states under the adjoint action of a generalization of the qubit $\pi/8$-gate. A specified stabilizer state maximally violates the inequality among all qudit states based on its Wigner negativity. The Bell operator not only serves as a measure for the singlet fraction but also quantifies the volume of Wigner negativity. Additionally, we show how a bipartite entangled qudit state can serve as a witness for contextuality when it exhibits Wigner negativity. Furthermore, we give deterministic Bell violations, as well as violations with a constant number of measurements, for the Bell state relying on operators innate to higher-dimensional systems than the qudit at hand.
@@
which straightforwardly extends to any number of qudits, in particular $\chi_{u,v}[\rho]=\mathrm{Tr}(\rho\,T_{-u}\otimes T_{-v})$ if $\rho$ is a two-qudit state. The discrete Fourier transform of the characteristic function is Gross’s Wigner function, a quasi-probability distribution that is real-valued and normalized, but can have negative values Gross2006 . Writing $u=(u_{x},u_{z})\in\mathbb{Z}_{d}^{2}$ and using the symplectic form $[u,v]:=u_{z}v_{x}-u_{x}v_{z}$, it is
@@
For a system of $n$ qudits, we denote $\bm{u}=({\bm{u}_{x}},{\bm{u}_{z}})\in\mathbb{Z}_{d}^{2n}$, and the operator $T_{\bm{u}}=\bigotimes_{i=1}^{n}T_{((u_{i})_{x},(u_{i})_{z})}$, similarly for $A_{\bm{u}}$, which also define analogously the Wigner function $W_{\bm{u}}[\rho]$ and characteristic function $\chi_{\bm{u}}[\rho]$ for a state $\rho$ on multiple qudits. In this context, we also use the inner product ${\bm{u}}\cdot{\bm{v}}=\sum_{i=1}^{n}u_{i}v_{i}$, which also extends to the symplectic form $[\bm{u},\bm{v}]=\bm{u}_{z}\bm{v}_{x}-\bm{u}_{x}\bm{v}_{z}$.
@@
A natural generalization of this inequality to qudits considers the $d^{\text{th}}$ root of unity $\omega=\exp{(2\pi\mathrm{i}/d)}$, instead of the second root of unity $-1$, such that
@@
$\langle\mathcal{B}[\rho]\rangle_{\mathrm{lhv}}\leq\max_{\alpha,\beta\in\mathbb{Z}_{d}^{d+1}}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}\chi_{u_{1},u_{2}}[\rho]\,\omega^{\alpha_{r(u_{1})}+\beta_{r(u_{2})}}\equiv B^{\max}_{\mathrm{lhv}}\,.$ (12)
@@
$B^{\max}_{\mathrm{lhv}}=\max_{\alpha,\beta\in\mathbb{Z}_{d}^{d+1}}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}W_{u_{1},u_{2}}[\rho]\,F(\alpha)_{u_{1}}F(\beta)_{u_{1}}\,.$ (13)
with the discrete Fourier transform of $\omega^{a_{r(u)}}$, $F(a)_{u}=\sum_{v\in\mathbb{Z}_{d}^{2}}\omega^{[u,v]+\alpha_{r(v)}}/d$. The values of $F(\alpha)_{u}$ range between $-1,0,\dots,d$, the sum of entries is $\sum_{u\in\mathbb{Z}_{d}^{2}}F(\alpha)_{u}=d$, and the vector’s norm is equally $\|F(\alpha)\|_{2}=d$. For an $\alpha\in\mathbb{Z}_{d}^{d+1}$ such that $F(\alpha)_{v}=d$ for some $v\in\mathbb{Z}_{d}^{2}$, it is $\alpha_{r(u)}=[a,u]$ for some $a\in\mathbb{Z}_{d}^{2}$, and $F(\alpha)_{w}=0$ for all $w\in\mathbb{Z}_{d}^{2}\setminus\{v\}$, which is precisely, the non-contextual value assignment used for the Wigner negativity witness or non-contextual inequality (11). To evaluate the Bell bound (13) further, the Cauchy-Schwarz inequality gives $B^{\max}_{\mathrm{lhv}}[\rho]\leq\|W[\rho]\|_{2}\|F(\alpha)\|_{2}\|F(\beta)\|_{2}\leq d^{2}\|W[\rho]\|_{2}$. To observe a Bell violation, it is necessary that $B_{\mathrm{lhv}}[\rho]<1$, which requires at least $\|W[\rho]\|_{2}\leq 1$. Since the non-contextual value assignments are a special case of the deterministic lhv models’, it is $B^{\max}_{\mathrm{nc}}\leq B^{\max}_{\mathrm{lhv}}$ such that Wigner negativity of the pure resource state $\rho$ as well as the state used for the quantum realization is again essential for nonlocality. Whether the state defines a notable Bell inequality beyond these criteria requires an analysis of the state’s characteristic or Wigner function. Bell states rotated by local unitary non-Clifford operators are prime candidates to define a Bell inequality in terms of their characteristic function.
@@
$\mathcal{B}_{\nu}^{\prime}=\sum_{x,z,t\in\mathbb{Z}_{d}}\chi^{\nu}_{(x,z),(x,-z)}\phantom{\cdot}U_{\nu}T_{(x,z)}U^{\dagger}_{\nu}\otimes T_{(x,t-z)}\,.$ (20)
@@
$2^{\ast\ast}$ $0.213$ $0.354$ $-0.707$ $-\phantom{1}0.707$ $0.213$
[…]
5 $0.724$ $0.447$ $-2.236$ $-\phantom{1}2.236$ $0.877$
7 $0.677$ $0.725$ $-4.406$ $-\phantom{1}4.406$ $0.824$
[…]
13 $0.442$ $1.102$ $-6.953$ $-10.651$ $0.616$
17 $0.437$ $1.251$ $-7.030$ $-14.728$
19 $0.449$ $1.437$ $-6.438$ $-16.755$
23 $0.371$ $1.531$ $-8.654$ $-20.795$
Table 1: Extremal values of the Wigner function $W^{\nu}_{v_{1},v_{2}}$ in Eq. (18) and Eq. (22), maximal values of its negativity volume $N^{\nu}:=N[\Phi_{\nu}]$ and the nonlocal bound $B^{\max}_{\mathrm{lhv}}$ from Eq. (12), all found through an exhaustive search, as well as values of $C_{\min}$ (Eq. (26)) for small prime dimension $d$. For $d\leq 7$, we saturate the bound of Eq. (27) and achieve $C_{\min}=d^{3}\min W^{\nu}_{v_{1},v_{2}}$. * For $d=3$, we instead use $\chi(f)=1+2\cos\left(8\pi/9\right)$ as we allow for cube roots of the character. ** For $d=2$, we use the Wigner function adapted to qubits for the Bell state and the $\pi/8$-gate in Eq. (33). Its minimum value is $-1/\sqrt{2}\approx-0.707$, corresponding to Tsirelson’s bound Cirelson1980 . Anyway, the Bell operator (9) and analysis thereof is not apt to qubits.
@@
As mentioned before, the case of $d=3$ differs from other odd prime dimensions in the sense that the approach with unitary cube operators does not work since all unitary operators defined by character polynomials are Clifford operators, since for $x\in\mathbb{Z}_{d}$ it is $x^{d}\sim x$. Therefore, Howard and Vala Howard2012 use the third root of the characters, $(-1)^{1/9}=\omega^{1/3}$ by defining $\nu_{k}=(6zk^{2}+2\gamma k+3k\epsilon)/3$ in Eq. (17). The Wigner function of the rotated Bell state is
@@
In fact, operators with a spectrum beyond the qudit characters can achieve stronger Bell violations for any odd prime dimension $d$. For example, the diagonal unitary operators $V_{{q}}=\sum_{k\in\mathbb{Z}_{d}}\omega^{k{q}}|k\rangle\langle k|$ where ${q}$ is a non-integer rational number, rotate the Pauli operators $X_{(q)}=V^{\phantom{\dagger}}_{{q}}XV^{\dagger}_{{q}}$, which, for ${q}=1/2$, lead to the Bell operator
$\mathcal{B}_{(1/2)}$ $=\frac{1}{d}\sum_{k\in\mathbb{Z}_{d}}X^{k}_{(1/2)}\otimes\left(X^{k}_{(-1/2)}+\omega^{-k}X^{k}_{(1/2)}\right)$
$+X^{k}_{(-1/2)}\otimes\left(X^{k}_{(1/2)}+\omega^{k}X^{k}_{(-1/2)}\right)\,.$ (24)
The corresponding Bell inequality is $\langle\mathcal{B}_{(1/2)}\rangle_{\,\mathrm{lhv}}\leq 3$. In contrast, the Bell state has an expectation value $\langle\Phi|\mathcal{B}_{(1/2)}|\Phi\rangle=4$ achieving a deterministic violation. One can even reduce the number of operators and only consider $k=1,d-1$ in Eq. (24). Then, an lhv model can achieve $\langle\mathcal{B}^{\prime}_{(1/2)}\rangle_{\,\mathrm{lhv}}\leq 3+\cos(4\pi/d)<4$. As a trade-off, the separation between the classical and quantum models grows smaller with increasing $d$.
@@
We have constructed several families of qudit Bell inequalities for any odd prime dimension. First, we propose a novel form of general Bell inequalities where a bipartite entangled state exhibiting Wigner negativity acts as a witness for nonlocality. Similarly to the $\pi/8$-rotation leading to the CHSH inequality, our Bell operator consists of stabilizer elements under local rotation of a generalization of the qubit $\pi/8$-gate characterized in Howard2012 . We illustrate how the Bell inequalities can be regarded as a natural extension of the CHSH inequality. The constituting measurements test correlations related to the rotated states’ Wigner negativity, a witness of non-classicality. Our family of Bell operators, as well as its quantum and classical bound, is easy to analyze and can be adapted to any entangled stabilizer state, which also reduces the number of measurements. The measurements can be generalized to any non-Clifford unitary operator, and we specifically discuss those from higher-degree polynomials over finite fields, which we find to achieve larger Wigner negativity and thus increase the separation between quantum and classical models. Moreover, nonlocal correlations can emerge with qudit Pauli operators when rotated by unitary operators with a non-character spectrum in bipartite systems. These operators have an obvious advantage, which is resorting to phases $\omega^{q}$ that are beyond the description of any local value assignment that can only resort to characters $\omega^{k}$. Such operations might arise from interactions with higher-dimensional systems, but the feasibility of the implementation of any non-Clifford operator present in this work remains an open question. Its study heavily depends on the platform of experimental realization with possible candidates being super-conducting circuits, cold atoms and optical systems involving GKP (Gottesman–Kitaev–Preskill) states PhysRevA.64.012310 .
@@ Appendix A Wigner Negativity
$=\frac{1}{d^{5}}\sum_{\begin{subarray}{c}j,k\in\mathbb{Z}_{d}\\ v_{1},v_{2}\in\mathbb{Z}_{d}^{2}\end{subarray}}\omega^{[u_{1},v_{1}]+[u_{2},v_{2}]}\langle jj|U^{\phantom{\dagger}}_{{f}}\,T_{v_{1}}U^{\dagger}_{{f}}\otimes T_{v_{2}}|kk\rangle$
$=\frac{1}{d^{5}}\sum_{\begin{subarray}{c}j,k\in\mathbb{Z}_{d}\\ v_{1},v_{2}\in\mathbb{Z}_{d}^{2}\end{subarray}}\omega^{[u_{1},v_{1}]+[u_{2},v_{2}]+k((v_{1})_{z}+(v_{2})_{z})+2^{-1}((v_{1})_{x}(v_{1})_{z}+(v_{2})_{x}(v_{2})_{z})}\langle jj|U^{\phantom{\dagger}}_{{f}}\,X^{(v_{1})_{x}}U^{\dagger}_{{f}}\otimes X^{(v_{2})_{x}}|kk\rangle$
$=\frac{1}{d^{5}}\sum_{\begin{subarray}{c}k\in\mathbb{Z}_{d}\\ v_{1},v_{2}\in\mathbb{Z}_{d}^{2}\end{subarray}}\delta_{(v_{1})_{x},(v_{2})_{x}}\,\omega^{[u_{1},v_{1}]+[u_{2},v_{2}]+k((v_{1})_{z}+(v_{2})_{z})+2^{-1}((v_{1})_{x}(v_{1})_{z}+(v_{2})_{x}(v_{2})_{z})+{f}_{k+(v_{1})_{x}}-{f}_{k^{\phantom{(1)}}}}$
${}^{(v_{x}:=(v_{1})_{x})}$ $=\frac{1}{d^{5}}\sum_{\begin{subarray}{c}v_{x},k\in\mathbb{Z}_{d}\\ (v_{1})_{z},(v_{2})_{z}\in\mathbb{Z}_{d}\end{subarray}}\omega^{((u_{1})_{z}+(u_{2})_{z})v_{x}-(u_{1})_{x}(v_{1})_{z}-(u_{2})_{x}(v_{2})_{z}+k((v_{1})_{z}+(v_{2})_{z})+2^{-1}v_{x}((v_{1})_{z}+(v_{2})_{z})+{f}_{k+v_{x}}-{f}_{k}}$

### v5

@@ Abstract
Nonlocality is an essential concept that distinguishes quantum from classical models and has been extensively studied in systems of qubits. For higher-dimensional systems, certain results for their two-level counterpart, like Bell violations with stabilizer states and Clifford operators, do not generalize. On the other hand, similar to continuous variable systems, Wigner negativity is necessary for nonlocality in qudit systems. We propose a new generalization of the CHSH inequality for qudits by inquiring correlations related to the Wigner negativity of stabilizer states under the adjoint action of a generalization of the qubit $\pi/8$-gate. A specified stabilizer state maximally violates the inequality among all qudit states based on its Wigner negativity. The Bell operator not only serves as a measure for the singlet fraction but also quantifies the volume of Wigner negativity. Additionally, we show how a bipartite entangled qudit state can serve as a witness for contextuality when it exhibits Wigner negativity. Furthermore, we identify rational-phase diagonal unitaries as the key resource that exactly reproduce the CGLMP violation with the maximally entangled state through simple phase-difference alignment.
@@
which straightforwardly extends to any number of qudits, in particular $\chi_{u,v}[\rho]=\mathrm{Tr}(\rho\,T_{-u}\otimes T_{-v})$ if $\rho$ is a two-qudit state. The discrete Fourier transform of the characteristic function is Gross’s Wigner function, a quasi-probability distribution that is real-valued and normalized, but can have negative values Gross2006 . Writing $u=(u_{x},u_{z})\in\mathbb{Z}_{d}^{2}$ and using the symplectic form $u,v]:=u_{z}v_{x}-u_{x}v_{z}$, it is
@@
For a system of $n$ qudits, we denote $\bm{u}=({\bm{u}_{x}},{\bm{u}_{z}})\in\mathbb{Z}_{d}^{2n}$, and the operator $T_{\bm{u}}=\bigotimes_{i=1}^{n}T_{((u_{i})_{x},(u_{i})_{z})}$, similarly for $A_{\bm{u}}$, which also define analogously the Wigner function $W_{\bm{u}}[\rho]$ and characteristic function $\chi_{\bm{u}}[\rho]$ for a state $\rho$ on multiple qudits. In this context, we also use the inner product ${\bm{u}}\cdot{\bm{v}}=\sum_{i=1}^{n}u_{i}v_{i}$, which also extends to the symplectic form $\bm{u},\bm{v}]=\bm{u}_{z}\bm{v}_{x}-\bm{u}_{x}\bm{v}_{z}$.
@@
A natural generalization of this inequality to qudits considers the $d^{\text{th}}$ root of unity $\omega=\exp{(2\pi\mathrm{i}/d)}$, instead of the second root of unity 1-1, such that
@@
$\langle\mathcal{B}[\rho]\rangle_{\mathrm{lhv}}\leq\hskip-2.84544pt\max_{\alpha,\beta\in\mathbb{Z}_{d}^{d+1}}\hskip-2.84544pt\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}\chi_{u_{1},u_{2}}[\rho]\,\omega^{\alpha_{r(u_{1})}+\beta_{r(u_{2})}}\equiv B^{\max}_{\mathrm{lhv}}\,.$ (12)
@@
$B^{\max}_{\mathrm{lhv}}=\hskip-2.84544pt\max_{\alpha,\beta\in\mathbb{Z}_{d}^{d+1}}\hskip-2.84544pt\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}W_{u_{1},u_{2}}[\rho]\,F(\alpha)_{u_{1}}F(\beta)_{u_{1}}\,.$ (13)
with the discrete Fourier transform of $\omega^{a_{r(u)}}$, $F(a)_{u}=\sum_{v\in\mathbb{Z}_{d}^{2}}\omega^{[u,v]+\alpha_{r(v)}}/d$. The values of $F(\alpha)_{u}$ range between $1,0,\dots,d$, the sum of entries is $\sum_{u\in\mathbb{Z}_{d}^{2}}F(\alpha)_{u}=d$, and the vector’s norm is equally $\|F(\alpha)\|_{2}=d$. For an $\alpha\in\mathbb{Z}_{d}^{d+1}$ such that $F(\alpha)_{v}=d$ for some $v\in\mathbb{Z}_{d}^{2}$, it is $\alpha_{r(u)}=[a,u]$ for some $a\in\mathbb{Z}_{d}^{2}$, and $F(\alpha)_{w}=0$ for all $w\in\mathbb{Z}_{d}^{2}\setminus\{v\}$, which is precisely, the non-contextual value assignment used for the Wigner negativity witness or non-contextual inequality (11). To evaluate the Bell bound (13) further, the Cauchy-Schwarz inequality gives $B^{\max}_{\mathrm{lhv}}[\rho]\leq\|W[\rho]\|_{2}\|F(\alpha)\|_{2}\|F(\beta)\|_{2}\leq d^{2}\|W[\rho]\|_{2}$. To observe a Bell violation, it is necessary that $B_{\mathrm{lhv}}[\rho]<1$, which requires at least $\|W[\rho]\|_{2}\leq 1$. Since the non-contextual value assignments are a special case of the deterministic lhv models’, it is $B^{\max}_{\mathrm{nc}}\leq B^{\max}_{\mathrm{lhv}}$ such that Wigner negativity of the pure resource state $\rho$ as well as the state used for the quantum realization is again essential for nonlocality. Whether the state defines a notable Bell inequality beyond these criteria requires an analysis of the state’s characteristic or Wigner function. Bell states rotated by local unitary non-Clifford operators are prime candidates to define a Bell inequality in terms of their characteristic function.
@@
$\mathcal{B}_{\nu}^{\prime}=\hskip-5.69046pt\sum_{x,z,t\in\mathbb{Z}_{d}}\hskip-5.69046pt\chi^{\nu}_{(x,z),(x,-z)}\phantom{\cdot}U_{\nu}T_{(x,z)}U^{\dagger}_{\nu}\otimes T_{(x,t-z)}\,.$ (20)
@@
$2^{\ast\ast}$ $0.213$ $0.354$ 0.707-0.707 $\phantom{1}0.707$ $0.213$
$3^{\ast}$ $0.844$ $0.293$ 0.879-0.879 $\phantom{1}0.879$ $0.960$
5 $0.724$ $0.447$ 2.236-2.236 $\phantom{1}2.236$ $0.877$
7 $0.677$ $0.725$ 4.406-4.406 $\phantom{1}4.406$ $0.829$
11 $0.535$ $0.914$ 4.211-4.211 $\phantom{1}8.595$ $0.774$
13 $0.442$ $1.102$ 6.953-6.953 10.651-10.651 $0.817$
17 $0.437$ $1.251$ 7.030-7.030 14.728-14.728 $0.736$
19 $0.449$ $1.437$ 6.438-6.438 16.755-16.755 $0.740$
23 $0.371$ $1.531$ 8.654-8.654 20.795-20.795 $0.666$
Table 1: Extremal values of the Wigner function $W^{\nu}_{v_{1},v_{2}}$ in Eq. (18) and Eq. (22), maximal values of its negativity volume $N^{\nu}:=N[\Phi_{\nu}]$ and the nonlocal bound $B^{\max}_{\mathrm{lhv}}$ from Eq. (12), all found through an , as well as values of $C_{\min}$ (Eq. (30)) for small prime dimension $d$. For $d\leq 7$, we saturate the bound of Eq. (31) and achieve $C_{\min}=d^{3}\min W^{\nu}_{v_{1},v_{2}}$. * For $d=3$, we instead use $\chi(f)=1+2\cos\left(8\pi/9\right)$ as we allow for cube roots of the character. ** For $d=2$, we use the Wigner function adapted to qubits for the Bell state and the $\pi/8$-gate in Eq. (37). Its minimum value is $1/\sqrt{2}\approx-0.707$, corresponding to Tsirelson’s bound Cirelson1980 . Anyway, the Bell operator (9) and analysis thereof is not apt to qubits.
@@
The case of $d=3$ differs from other odd prime dimensions in the sense that the approach with unitary cube operators does not work since all unitary operators defined by character polynomials are Clifford operators, since for $x\in\mathbb{Z}_{d}$ it is $x^{d}\sim x$. Therefore, Howard and Vala Howard2012 use the third root of the characters, $(-1)^{1/9}=\omega^{1/3}$ by defining $\nu_{k}=(6zk^{2}+2\gamma k+3k\epsilon)/3$ in Eq. (17). The Wigner function of the rotated Bell state is
@@
Operators with a spectrum beyond the qudit characters can also achieve strong Bell violations using the maximally entangled Bell state for any odd prime dimension $d$. For example, consider the diagonal unitary operators
$V_{{q}}=\sum_{k\in\mathbb{Z}_{d}}\omega^{k\{q\}}|k\rangle\langle k|$ (24)
where $\{q\}$ is a non-integer rational number, and rotate the Pauli operators $X_{(q)}=V^{\phantom{\dagger}}_{{q}}XV^{\dagger}_{{q}}$.
The Hermitian Bell operator
[…]
for rational numbers $q_{0},q_{1},p_{0},p_{1}$ and weights $\gamma_{k}=\gamma_{-k}$ is the discrete Fourier transform of a Collins-Gisin-Linden-Massar-Popescu (CGLMP) functional Collins2001 ,
$I_{\mathrm{CGLMP}}$ $=\sum_{r=0}^{d-1}\,W_{r}\,\Big{(}P(r_{00}=r)+P(r_{01}=r)$
$\hskip 42.67912pt+P(r_{10}=r)-P(r_{11}=r)\Big{)}\,$, (26)
The CGLMP choice corresponds to $(q_{0},q_{1};p_{0},p_{1})=(0,1/2;1/4,-1/4)$ and to the symmetric $\gamma_{k}:=1-2\,\min\{k,d-k\}/(d-1)$, the Fourier transform of $W_{r}$. With these weights, the Bell operator $\mathcal{B}_{\gamma}$ satisfies the local bound
$\max\langle I_{\mathrm{CGLMP}}\rangle_{\mathrm{lhv}}\leq 2\,.$ (27)
Other even weights $\gamma_{k}$ also define valid Bell functionals; the ramp $\gamma_{k}$ is distinguished by a tight, closed-form lhv bound Collins2001 . For the quantum values, note that $\langle\Phi|(X_{(q)})^{k}\otimes(X_{(p)})^{k}\,|\Phi\rangle=\omega^{k(p-q)}$ such that the maximally entangled state achieves
$\langle\Phi|\mathcal{B}_{\gamma}|\Phi\rangle$ $=\sum_{k\in\mathbb{Z}_{d}^{\prime}}\gamma_{k}\Big{[}\omega^{k(p_{0}-q_{0})}+\omega^{k(p_{1}-q_{0})}$
$\hskip 42.67912pt+\omega^{k(p_{0}-q_{1})}-\omega^{k(p_{1}-q_{1})}\Big{]}\,.$ (28)
The freedom needed to generate the non-local violation with the operator (25) using the Bell state is in the phase differences pqp-q. Choosing the rational offsets $(0,\tfrac{1}{2};\tfrac{1}{4},-\tfrac{1}{4})$ aligns all modes constructively in (Non-character operators.) and lead to a violation of the Bell inequality (27Collins2001 . This shows that the non-integer phases supplied by $V_{q}$ control the phase of each Fourier mode $k$ via the differences pqp-q and achieve the quantum advantage over any local model. Taken together, these results show that diagonal unitaries with non-integer spectra $V_{q}$ supply the precise phase control needed to obtain a bipartite qutrit violation with non-character operators, and for any odd prime $d$, recover the CGLMP violation by aligning all Fourier modes via the phase differences.
@@
We have constructed several families of qudit Bell inequalities for any odd prime dimension. First, we propose a novel form of general Bell inequalities where a bipartite entangled state exhibiting Wigner negativity acts as a witness for nonlocality. Similarly to the $\pi/8$-rotation leading to the CHSH inequality, our Bell operator consists of stabilizer elements under local rotation of a generalization of the qubit $\pi/8$-gate characterized in Howard2012 . We illustrate how the Bell inequalities can be regarded as a natural extension of the CHSH inequality. The constituting measurements test correlations related to the rotated states’ Wigner negativity, a witness of non-classicality. Our family of Bell operators, as well as its quantum and classical bound, is easy to analyze and can be adapted to any entangled stabilizer state, which also reduces the number of measurements. The measurements can be generalized to any non-Clifford unitary operator, and we specifically discuss those from higher-degree polynomials over finite fields, which we find to achieve larger Wigner negativity and thus increase the separation between quantum and classical models.
Moreover, nonlocal correlations already arise when qudit Pauli measurements are conjugated by diagonal unitaries with rational, non-character eigenphases. These rotations supply phase offsets that coherently align the correlation modes for the maximally entangled state, while any deterministic local model - restricted to character phases - cannot reproduce this sensitivity, yielding the bipartite qutrit violation and, in higher odd prime dimensions, the standard CGLMP violation. Such operations might arise from interactions with higher-dimensional systems, but the feasibility of the implementation of any non-Clifford operator present in this work remains an open question. Its study heavily depends on the platform of experimental realization with possible candidates being super-conducting circuits, cold atoms and optical systems involving Gottesman-Kitaev-Preskill (GKP) states PhysRevA.64.012310 .
@@ Appendix A Wigner Negativity
$=\frac{1}{d^{5}}\hskip-5.69046pt\sum_{\begin{subarray}{c}j,k\in\mathbb{Z}_{d}\\ v_{1},v_{2}\in\mathbb{Z}_{d}^{2}\end{subarray}}\hskip-5.69046pt\omega^{[u_{1},v_{1}]+[u_{2},v_{2}]}\langle jj|U^{\phantom{\dagger}}_{{f}}\,T_{v_{1}}U^{\dagger}_{{f}}\otimes T_{v_{2}}|kk\rangle$
$=\frac{1}{d^{5}}\hskip-5.69046pt\sum_{\begin{subarray}{c}j,k\in\mathbb{Z}_{d}\\ v_{1},v_{2}\in\mathbb{Z}_{d}^{2}\end{subarray}}\hskip-5.69046pt\omega^{[u_{1},v_{1}]+[u_{2},v_{2}]+k((v_{1})_{z}+(v_{2})_{z})+2^{-1}((v_{1})_{x}(v_{1})_{z}+(v_{2})_{x}(v_{2})_{z})}\langle jj|U^{\phantom{\dagger}}_{{f}}\,X^{(v_{1})_{x}}U^{\dagger}_{{f}}\otimes X^{(v_{2})_{x}}|kk\rangle$
$=\frac{1}{d^{5}}\hskip-5.69046pt\sum_{\begin{subarray}{c}k\in\mathbb{Z}_{d}\\ v_{1},v_{2}\in\mathbb{Z}_{d}^{2}\end{subarray}}\hskip-5.69046pt\delta_{(v_{1})_{x},(v_{2})_{x}}\,\omega^{[u_{1},v_{1}]+[u_{2},v_{2}]+k((v_{1})_{z}+(v_{2})_{z})+2^{-1}((v_{1})_{x}(v_{1})_{z}+(v_{2})_{x}(v_{2})_{z})+{f}_{k+(v_{1})_{x}}-{f}_{k^{\phantom{(1)}}}}$
$^{(v_{x}:=(v_{1})_{x})}$ $=\frac{1}{d^{5}}\hskip-5.69046pt\sum_{\begin{subarray}{c}v_{x},k\in\mathbb{Z}_{d}\\ (v_{1})_{z},(v_{2})_{z}\in\mathbb{Z}_{d}\end{subarray}}\hskip-5.69046pt\omega^{((u_{1})_{z}+(u_{2})_{z})v_{x}-(u_{1})_{x}(v_{1})_{z}-(u_{2})_{x}(v_{2})_{z}+k((v_{1})_{z}+(v_{2})_{z})+2^{-1}v_{x}((v_{1})_{z}+(v_{2})_{z})+{f}_{k+v_{x}}-{f}_{k}}$